# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 3.4. $\ell$-integrality of Rankin–Selberg $L$-values
We first introduce normalised test vectors.
Let $\ell$ be a prime that is coprime to the primes in $S_{0}$. Fix embeddings $\iota_{\ell}:\overline{{\mathbb{Q}}}\hookrightarrow\overline{{\mathbb{Q}}}_{\ell}$ and $\iota:\overline{{\mathbb{Q}}}\hookrightarrow{\mathbb{C}}$. Let $L$ be a sufficiently large number field as in §2.1.5. Let $O_{\pi,\ell}\subset{\mathbb{C}}_{\ell}$ be the completion of the ring of integers of $L$ with respect to the prime ${\mathfrak{l}}|\ell$ of $L$ determined by the embedding $\iota_{\ell}$.
For an $\ell$-primitive $x\in\bigotimes_{\sigma\in\Sigma}V_{k_{\sigma}-2}(O_{\pi,\ell})$, let $\langle{\ ,\ }\rangle_{x}$ be an $\ell$-optimally normalized invariant pairing [^12] It is unique up to multiplication by elements of $O_{\pi,\ell}^{\times}$. on $\bigotimes_{\sigma\in\Sigma}V_{k_{\sigma}-2}(O_{\pi,\ell})$.
Recall that $x_{0}=\otimes_{\sigma\in\Sigma}x_{0_{\sigma}}\in\ \bigotimes_{\sigma\in\Sigma}V_{k_{\sigma}-2}(O_{\pi,\ell})$ where $x_{0_{\sigma}}=X^{\frac{k_{\sigma}-2}{2}}Y^{\frac{k_{\sigma}-2}{2}}$ is $\ell$-primitive. From now on, fix
@@
and let
$\varphi\in{\rm M}_{k}(\widehat{R}^{\times},L)[\pi_{f}]^{\prod_{v|\mathfrak{N}^{-},v\notin S_{0}}\iota_{v}K_{v}^{\times}}$
be a basis such that $f=\Psi(x_{0}\otimes\varphi)$ is the test vector as in Definition 3.8. We choose $\widehat{\varphi}$ to be $\ell$-optimal with respect to $x$ in the sense of Definition 2.2. Recall that $\varsigma_{\ell}\in{\mathrm{GL}}_{2}(O_{F_{\ell}})$ so that $\rho_{k,\ell}(\varsigma_{\ell})^{-1}x_{0}$ is $\ell$-primitive. Also note that if $\ell$ is coprime to $\mathfrak{N}^{+}$, then $\varsigma_{\ell}=1$. For brevity, the present choice of $f$ is said to be $\ell$-optimal.
@@
where the Petersson norms of $\phi_{\lambda}$ and $f$ are normalized as in §3.3.1.
@@
Let $\chi\in{\mathcal{X}}_{S_{0}}$ be with conductor $\mathfrak{c}$} such that $O_{K,{\mathfrak{c}}_{1}}^{\times}:O^{\times}]=1$. Then we have
@@
where $2^{r}:=\frac{\kappa_{{\mathfrak{c}}_{1}}^{2}}{2^{c}}$ as in Theorem 3.12.
@@
In view of the $\ell$-optimality of $f$, the toric period $P(\varsigma^{({\mathfrak{c}})},\varphi,\chi)$ is $\ell$-integral. Hence, the assertion follows from the explicit Waldspurger formula as in  Theorem 3.12. ∎
@@
If the norm of ${\mathfrak{c}}_{1}$ is sufficiently large, then $O_{K,{\mathfrak{c}}_{1}}^{\times}:O^{\times}]=1.$
@@ 4. Non-vanishing of Rankin–Selberg $L$-values modulo $\ell$: CM case
For an odd prime $\ell$ and a prime $\mathfrak{p}$} of a totally real number field $F$ coprime to $\ell$, this section presents mod $\ell$ non-vanishing of Rankin–Selberg $L$-values in $\mathfrak{p}$}-anticyclotic twist family of a CM Hilbert newform over $F$. We first consider the case $\deg{\mathfrak{p}}=1$, and the main results are Theorems 4.14 and 4.17 if $\mathfrak{p}$} is inert or split in the CM field, respectively. If $\deg{\mathfrak{p}}>1$, then the main result is Theorem 4.31. The latter case turns out to be more involved: we partition the $\mathfrak{p}$}-anticyclotomic family into infinitely many subfamilies indexed by ${\mathbb{Z}}_{p}$-quotients of the underlying Galois group $\Gamma_{\mathfrak{p}}\simeq{\mathbb{Z}}_{p}^{\deg{\mathfrak{p}}}$ and establish uniform non-vanishing over each subfamily.
@@ 4.1.1. Equidistribution of special points
We describe a special case of the main result of [12], which is based on Ratner’s ergodicity of unipotent flows.
@@
of a totally imaginary quadratic field extension $K$ of $F$. Let $F_{+}\subset F^{\times}$ be the subgroup of totally positive elements. Write $\overline{K}^{\times}$, $\overline{B}^{\times}$, $\overline{F}_{+}^{\times}$ for the closure of $K^{\times}$, $B^{\times}$, $F_{+}^{\times}$ in $\widehat{K}^{\times}$, $\widehat{B}^{\times}$, $\widehat{F}^{\times}$ respectively.
@@
The group $\widehat{B}^{\times}$ acts on via right multiplication on ${\rm CM}_{K}$ and $X$, and via multiplication by the reduced norm on the third space. Similarly, there is a left action of $\widehat{K}^{\times}$ on ${\rm CM}_{K}$ and $Z$. Let $i:{\rm{CM}}_{K}\to X$ be the natural quotient map and $c:X\to Z$ the map induced by the reduced norm.
@@
The following key result is a special case of [12, Cor. 2.10].
@@
Let $D_{0}$ be a set of representatives of $\Delta^{\mathrm{alg}}$ in $\widehat{F}^{\times}\prod_{v|D_{K/F},v\nmid{\mathfrak{p}}}K_{v}^{\times}\prod_{v|{\mathfrak{c}}}O_{K_{v}}^{\times}$, and $D_{1}$ that of $\Delta/\Delta^{\mathrm{alg}}$ in $\widehat{K}^{\times}$. Then $D:=D_{1}D_{0}$ is a set of representatives of $\Delta$ in $\widehat{K}^{\times}$.
We recall the following (cf. [13, Sec. 2.2]).
@@
The elements in $D_{1}$ are pairwise distinct modulo image of $\widehat{K}^{({\mathfrak{p}}),\times}\cdot(K^{\times}F_{\mathfrak{p}}^{\times})$ in ${\rm{G}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}$.
@@ 4.1.2. Non-Eisenstein functions: generalities
For $f\in{\rm M}_{2}(U,A)$, write $f=f^{+}+f^{-}$ with $f^{\varepsilon}$ supported on $X_{U}^{\varepsilon}$.
@@ 4.1.3. An independence of CM values
Let $\mathbb{F}$} be a finite extension of $\mathbb{F}$$_{\ell}$. Let $(\beta_{\tau})_{\tau\in D_{1}}$ be a sequence of elements in $\mathbb{F}$} such that $\beta_{\tau_{1}}\neq 0$ for some $\tau_{1}$. Let $f\in{\rm M}_{2}(U,{\mathbb{F}})$ and suppose that ${\rm N}(U)\subset{\rm N}(\widehat{K}^{\times})$. Let $\mathfrak{c}$} be an integral ideal of $F$ coprime to ${\mathfrak{p}}D_{K/F}D_{B}$.
@@
    Suppose that $\mathfrak{p}$} is inert in $K$ and $f=f^{\varepsilon}$ is non-Eisenstein. Then there exists an integer $n_{0}$ such that for any $n>n_{0}$ of the fixed parity determined by $(-1)^{n}\eta_{K/F}({\mathfrak{c}})\equiv\varepsilon\pmod{2}$, we have
@@
    Suppose that $\mathfrak{p}$} splits in $K$, and that $f=f^{{\mathrm{sgn}}(\eta_{K/F}({\mathfrak{c}}))}$ is non-Eisenstein. Then there exists an integer $n_{0}$ such that for any $n>n_{0}$, we have
@@
For integers $n\geq 1$, put
$P_{n}:=P_{0}\delta_{n}\in{\mathcal{H}},$ (4.1)
@@
    We have $y,z\in X_{U}^{\varepsilon}$. For $n\geq 1$, note that $c(i(P_{n}))=c(i(P_{n+2}))\pmod{{\rm{N}}(\widehat{K}^{\times})}$, and
@@
    and $f(y)\neq f(z)$. Then the assertion follows just as in part (a).
@@
Our non-vanishing problems concern toric periods of a Hilbert modular CM form of arbitrary weight. Via congruences, the non-vanishing is closely related to that of a parallel weight two modular form as in the proposition.
@@ 4.1.4. Non-Eisenstein functions: under Hecke action
Let $U\subset\widehat{B}^{\times}$ be an open compact subgroup of the form $\prod_{v}U_{v}$.
@@
Let $v$ be a finite place such that $B_{v}$ splits and $U_{v}=U_{0}({\mathfrak{p}}_{v}^{k})_{v}$ for some $k$. For $\beta_{1},\cdots,\beta_{s}\in A$, define ${\mathcal{R}}\in{\mathrm{End}}({\rm M}_{2}(U,A))$ by
@@
Then ${\mathcal{R}}:{\rm S}_{2}(U,A)\to{\rm S}_{2}(A)$ is an injective map (cf. [11, Lem. 5.5]).
Let $A$ be the ring of integers of a finite extension of $\mathbb{Q}$$_{\ell}$ and $\varpi$ a uniformiser of $A$.
@@
Let $R\subset B$ be an $O$-order. Let $\widehat{\varphi}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ be the $\ell$-adic avatar of an eigenform $\varphi$ with level $\widehat{R}^{\times}$. Let $\mathfrak{q}$$\nmid\ell$ be a prime of $F$ non-split in $K$ and $q$ the cardinality of the residue field of $F_{\mathfrak{q}}$. Suppose that $\ell\nmid q(q^{2}-1)$, and that $\varphi_{\mathfrak{q}}$ is the newform [^13] That is the representation generated by $f_{\mathfrak{q}}$ is an irreducible representation of $B_{\mathfrak{q}}^{\times}$ for which the $R_{\mathfrak{q}}^{\times}$-invariant subspace is of dimensional $1$. with level $R_{\mathfrak{q}}^{\times}$. Let $\chi$ be a character of $K_{\mathfrak{q}}^{\times}$ valued in $A$. Suppose that
$\gamma:=\frac{1}{{\mathrm{vol}}(K_{\mathfrak{q}}^{\times}/F_{\mathfrak{q}}^{\times})}\int_{K_{\mathfrak{q}}^{\times}/F_{\mathfrak{q}}^{\times}}\frac{(\pi(t)\varphi_{\mathfrak{q}},\varphi_{\mathfrak{q}})_{\mathfrak{q}}}{(\varphi_{\mathfrak{q}},\varphi_{\mathfrak{q}})_{\mathfrak{q}}}\chi(t)d^{\times}t$
is an $\ell$-adic unit, where $(\ ,\ )_{\mathfrak{q}}$ is a non-degenerate $B_{\mathfrak{q}}^{\times}$-invariant Hermitian pairing on $\pi$. Moreover, suppose that $\widehat{\varphi}$ is $\ell$-optimal [^14] That is, $\langle{\widehat{\varphi},x}\rangle_{x}\pmod{\varpi}$ is a well defined non-zero weight two form. with respect to some $x\in P_{k-2\Sigma}(A)$. Put
$H=\sum_{\iota\in K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})}\chi(t)\pi(t)\widehat{\varphi}$
@@
$H^{\prime}=\sum_{g\in R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}}\pi(g)H\in\pi^{\widehat{R}^{\times}}.$
If $\langle{H^{\prime},x}\rangle_{x}$ is non-zero modulo $\varpi$, then so is $\langle{H,x}\rangle_{x}$. In the following we establish the former non-vanishing.
Let $\langle{\ ,\ }\rangle$ be the $G({\mathbb{A}}_{f}^{(\ell)})$-invariant pairing on ${\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ given by
$\langle{f,h}\rangle=\sum_{[g]\in B^{\times}\widehat{F}^{\times}\backslash\widehat{B}^{\times}/\widehat{R}^{\times}}\frac{\langle{f(g),h(g)}\rangle_{x}}{w_{g}},$
@@
$w_{g}\in A^{\times}$
since $R_{g}^{1}:=\{x\in R_{g}^{\times}\ |\ {\rm N}(x)=1\}\subset(gR_{\mathfrak{q}}^{\times}g^{-1})_{\text{tor}}$ and the prime factors of $R_{g}^{1}$ divide $q(q^{2}-1)$.
Note that $\langle{H^{\prime},x}\rangle_{x}$ is non-zero modulo $\varpi$ if and only if there exists $H^{\prime\prime}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ with $H^{\prime\prime}(\widehat{B}^{\times})\subset A\rho(R_{\ell}^{\times})x$ such that
$v_{\ell}(\langle{H^{\prime},H^{\prime\prime}}\rangle)=0.$
@@
$\langle{H^{\prime},H^{\prime\prime}}\rangle=$ $\langle{\sum_{g\in R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}}\pi(g)H,H^{\prime\prime}}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\langle{H,H^{\prime\prime}}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\langle{\sum_{t\in K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})}\chi(t)\pi(t)\widehat{\varphi},H^{\prime\prime}}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\cdot\gamma\#K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})\cdot\langle{\widehat{\varphi},H^{\prime\prime}}\rangle.$
Here the last equality follows from the multiplicity one of invariant pairings on $\pi$ (cf. [33, Lem. 2.6]) and the definition of $\gamma$.
Note that $R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}$, $\gamma$ and $K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})$ are $\ell$-adic units. Since $\langle{\widehat{\varphi},x}\rangle_{x}$ is non-zero modulo $\varpi$, the existence of $H^{\prime\prime}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ such that $H^{\prime\prime}(\widehat{B}^{\times})\subset A\rho(R_{\ell}^{\times})x$ and $\langle{\widehat{\varphi},H^{\prime\prime}}\rangle$ is non-zero modulo $\varpi$ follows, concluding the proof.

### v2

@@ 3.4. $\ell$-integrality of Rankin–Selberg $L$-values
In this subsection we present an $\ell$-integrality of Rankin–Selberg $L$-values in anticyclotomic twist family, based on the Waldspurger formula in §3.3.2 and $\ell$-integrality of quaternionic modular forms in §2.3.
Let the setting be as in §3.1.1. We first introduce normalised test vectors.
Let $\ell$ be a prime that is coprime to the primes in $S_{0}$. Fix embeddings $\iota_{\ell}:\overline{{\mathbb{Q}}}\hookrightarrow\overline{{\mathbb{Q}}}_{\ell}$ and $\iota:\overline{{\mathbb{Q}}}\hookrightarrow{\mathbb{C}}$. Let $L$ be a sufficiently large number field as in §2.3.3 so that it contains the Hecke field of $\pi$. Let $O_{\pi,\ell}\subset{\mathbb{C}}_{\ell}$ be the completion of the ring of integers of $L$ with respect to the prime ${\mathfrak{l}}|\ell$ of $L$ determined by the embedding $\iota_{\ell}$.
Recall that $x_{0}=\otimes_{\sigma\in\Sigma}x_{0_{\sigma}}\in\ \bigotimes_{\sigma\in\Sigma}V_{k_{\sigma}-2}(O_{\pi,\ell})$ where $x_{0_{\sigma}}=X^{\frac{k_{\sigma}-2}{2}}Y^{\frac{k_{\sigma}-2}{2}}$ is $\ell$-primitive and $K_{\sigma}^{\times}$-invariant (cf. §2.3.1). From now on, fix
@@
with $\varsigma_{\ell}$ as in §2.4.1, let $\pi_{f}$ be the finite part of $\pi$ and let
$\varphi\in{\rm M}_{k}(\widehat{R}^{\times},L)[\pi_{f}]^{\prod_{v|\mathfrak{N}^{-},v\notin S_{0}}{\iota_{\varsigma_{v}}K_{v}^{\times}}}$
be a basis such that $f=\Psi(x_{0}\otimes\varphi)$ is the test vector as in Definition 3.8. We choose $\widehat{\varphi}$ to be $\ell$-optimal with respect to $x$ in the sense of Definition 2.2.
For brevity, the present choice of $f$ is said to be $\ell$-optimal.
@@
where the Petersson norms of $\phi$ and $f$ are normalized as in §3.3.1.
@@
Let $(\pi,\chi)$ be a self-dual pair as in §3.1. Let $\ell$ be a rational prime that is coprime to the primes in $S_{0}$. Then for any $\chi\in{\mathcal{X}}_{S_{0}}$ with conductor $\mathfrak{r}$} such that $O_{K,{\mathfrak{r}}_{1}}^{\times}:O^{\times}]=1$ for ${\mathfrak{r}}_{1}=\prod_{\begin{subarray}{c}v\in S_{0}\\ \text{non-split}\end{subarray}}{\mathfrak{p}}_{v}^{r_{v}}$, we have
@@
where $2^{r}:=\frac{\kappa_{{\mathfrak{r}}_{1}}^{2}}{2^{c}}$ and $\Sigma_{0}$ is as in Theorem 3.14.
@@
Since $x_{0}$ is $K_{\infty}^{\times}$-invariant, we have
$\langle{x_{0},\varphi(x_{\mathfrak{r}}(a))}\rangle_{k-2\Sigma}=\langle{x,\widehat{\varphi}(x_{\mathfrak{r}}(a))}\rangle_{k-2\Sigma}.$
So in view of the $\ell$-optimality of $f$, the toric period $P(\varsigma^{({\mathfrak{r}})},\varphi,\chi)$ defined in §3.3.1 is $\ell$-integral. Hence, the assertion follows from the explicit Waldspurger formula as in  Theorem 3.14. ∎
@@
If the norm of ${\mathfrak{r}}_{1}$ is sufficiently large, then $O_{K,{\mathfrak{r}}_{1}}^{\times}:O^{\times}]=1.$
@@ 4. Non-vanishing of Rankin–Selberg $L$-values modulo $\ell$: CM case
For an odd prime $\ell$, a prime $\mathfrak{p}$} of a totally real number field $F$ coprime to $\ell$ and a CM quadratic extension $K$ of $F$, this section presents mod $\ell$ non-vanishing of Rankin–Selberg $L$-values in the $\mathfrak{p}$}-anticyclotic twist family of a CM Hilbert newform over $F$. The main results are Theorems 4.15 and 4.19 if $\mathfrak{p}$} is inert or split in $K$ respectively. We refer to  §1.4.3 for an outline of the strategy.
@@ 4.1.1. Equidistribution of special points
We describe an equidistribution of special points on a Shimura set based on Ratner’s ergodicity of unipotent flows (cf. [11]).
@@
of a totally imaginary quadratic field extension $K$ of $F$. Let $F_{+}^{\times}\subset F^{\times}$ be the subgroup of totally positive elements. Write $\overline{K}^{\times}$, $\overline{B}^{\times}$, $\overline{F}_{+}^{\times}$ for the closure of $K^{\times}$, $B^{\times}$, $F_{+}^{\times}$ in $\widehat{K}^{\times}$, $\widehat{B}^{\times}$, $\widehat{F}^{\times}$ respectively.
@@
The group $\widehat{B}^{\times}$ acts on via right multiplication on ${\rm CM}_{K}$ and $X$, and via multiplication by the reduced norm $\rm{N}$ on the third space. Similarly, there is a left action of $\widehat{K}^{\times}$ on ${\rm CM}_{K}$ and $Z$. Let $i:{\rm{CM}}_{K}\to X$ be the natural map and $c:X\to Z$ the map induced by the reduced norm.
@@
The following key result is a special case of [11, Cor. 2.10].
@@
In the following we will introduce an explicit $\mathcal{S}$} for our later application.
@@
Let $D_{1}$ be a set of representatives of $\Delta/\Delta^{\mathrm{alg}}$ in $\widehat{K}^{\times}$.
We recall the following (cf. [12, Sec. 2.2]).
@@
The elements in $D_{1}$ are pairwise distinct modulo image of $\widehat{K}^{({\mathfrak{p}}),\times}\cdot(K^{\times}F_{\mathfrak{p}}^{\times})$ in ${\rm{P}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}$.
@@ 4.1.2. Non-Eisenstein functions: generalities
For $f\in{\rm M}_{2}(U,A)$, write
$f=f^{+}+f^{-}$ (4.1)
with $f^{\varepsilon}$ supported on $X_{U}^{\varepsilon}$.
Remark 4.4.
For a higher weight modular form $f\in{\rm M}_{k}(U,A)$, we define a decomposition $f=f^{+}+f^{-}$ analogous to (4.1).
@@ 4.1.3. An independence of CM values
Let $\mathbb{F}$} be a finite extension of $\mathbb{F}$$_{\ell}$. Let $(\beta_{\tau})_{\tau\in D_{1}}$ be a sequence of elements in $\mathbb{F}$} such that $\beta_{\tau_{1}}\neq 0$ for some $\tau_{1}$. Let $f\in{\rm M}_{2}(U,{\mathbb{F}})$ and suppose that ${\rm N}(U)\subset F_{+}^{\times}{\rm N}(\widehat{K}^{\times})$. Let $\mathfrak{c}$} be an integral ideal of $F$ coprime to ${\mathfrak{p}}D_{K/F}D_{B}$.
@@
    Suppose that $\mathfrak{p}$} is inert in $K$ and $f=f^{\varepsilon}$ is non-Eisenstein. Then there exists an integer $n_{0}$ such that for any $n>n_{0}$ of the fixed parity determined by $\varepsilon=(-1)^{n}\eta_{K/F}({\mathfrak{c}})$, we have
@@
    Suppose that $\mathfrak{p}$} is split in $K$, and that $f=f^{{\mathrm{sgn}}(\eta_{K/F}({\mathfrak{c}}))}$ is non-Eisenstein. Then there exists an integer $n_{0}$ such that for any $n>n_{0}$, we have
@@
For $n\geq 0$, let $\delta_{n}\in{\mathrm{GL}}_{2}(F_{\mathfrak{p}})$ be the family of elements as in §2.4.1 and put
$P_{n}:=P_{0}\delta_{n}\in{\mathcal{H}}.$ (4.2)
@@
    We have $y,z\in X_{U}^{\varepsilon}$. Note that $c(i(P_{n}))=c(i(P_{n+2}))\pmod{{\rm{N}}(\widehat{K}^{\times})}$, and
@@
    and $f(y)\neq f(z)$. Then the assertion follows just as in part a).
@@
Our non-vanishing problems concern toric periods of a CM Hilbert modular form of arbitrary weight. Via congruences, the non-vanishing is closely related to that of a parallel weight two modular form as in the proposition.
@@ 4.1.4. Non-Eisenstein functions: under Hecke action
Let $U\subset\widehat{B}^{\times}$ be an open compact subgroup of the form $\prod_{v\nmid\infty}U_{v}$. Let $A$ be a commutative ring.
@@
Let $v$ be a finite place such that $B_{v}$ splits and $U_{v}=U_{0}({\mathfrak{p}}_{v}^{k})_{v}$ for some $k$. For $\beta_{1},\cdots,\beta_{s}\in A$, define ${\mathcal{R}}\in{\mathrm{End}}({\rm M}_{2}(A))$ by
@@
Then ${\mathcal{R}}:{\rm S}_{2}(U,A)\to{\rm S}_{2}(A)$ is an injective map (cf. [10, Lem. 5.5]).
Let $A$ be the ring of integers of a finite extension of $\mathbb{Q}$$_{\ell}$ and $\varpi$ a uniformiser of $A$. We have the following key mod $\ell$ refinement of Lemma 3.12.
@@
Let $R\subset B$ be an $O$-order. Let $\widehat{\varphi}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ be the $\ell$-adic avatar of an eigenform $\varphi$ with level $\widehat{R}^{\times}$. Let $K\subset B$ be an $F$-subalgebra, which is a totally imaginary quadratic extension of $F$ and $\mathfrak{q}$} be a prime of $F$ non-split in $K$ and $q$ the cardinality of the residue field of $F_{\mathfrak{q}}$. Suppose that $\ell\nmid q(q^{2}-1)$, and that $\varphi=\varphi_{\mathfrak{q}}\otimes\varphi^{({\mathfrak{q}})}$ with $\varphi_{\mathfrak{q}}$ is the newform [^18] That is the representation $\pi_{\mathfrak{q}}$ generated by $\varphi_{\mathfrak{q}}$ is an irreducible representation of $B_{\mathfrak{q}}^{\times}$ for which the $R_{\mathfrak{q}}^{\times}$-invariant subspace is of dimensional $1$. with level $R_{\mathfrak{q}}^{\times}$. Let $\chi_{\mathfrak{q}}$ be a character of $K_{\mathfrak{q}}^{\times}/F_{\mathfrak{q}}^{\times}$ valued in $A$. Suppose that
$\gamma:=\frac{1}{{\mathrm{vol}}(K_{\mathfrak{q}}^{\times}/F_{\mathfrak{q}}^{\times})}\int_{K_{\mathfrak{q}}^{\times}/F_{\mathfrak{q}}^{\times}}\frac{(\pi_{\mathfrak{q}}(t)\varphi_{\mathfrak{q}},\varphi_{\mathfrak{q}})_{\mathfrak{q}}}{(\varphi_{\mathfrak{q}},\varphi_{\mathfrak{q}})_{\mathfrak{q}}}\chi_{\mathfrak{q}}(t)d^{\times}t$
is an $\ell$-adic unit, where $(\ ,\ )_{\mathfrak{q}}$ is a non-degenerate $B_{\mathfrak{q}}^{\times}$-invariant pairing on $\pi_{\mathfrak{q}}$. Moreover, suppose that $\widehat{\varphi}$ is $\ell$-optimal [^19] That is, $\langle{\widehat{\varphi},x}\rangle_{k-2\Sigma}\pmod{\varpi}$ is a well defined non-zero weight two form. with respect to some primitive $x\in P_{k-2\Sigma}(A)$. Put
$H=\sum_{\iota\in K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})}\chi_{\mathfrak{q}}(t)\pi_{\mathfrak{q}}(t)\widehat{\varphi}$
@@
$H^{\prime}=\sum_{g\in R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}}\pi_{\mathfrak{q}}(g)H\in\pi^{\widehat{R}^{\times}}.$
If $\langle{x,H^{\prime}}\rangle_{k-2\Sigma}$ is non-zero modulo $\varpi$, then so is $\langle{x,H}\rangle_{k-2\Sigma}$. In the following we establish the former non-vanishing.
[…]
$\langle{f,h}\rangle=\sum_{[g]\in B^{\times}\widehat{F}^{\times}\backslash\widehat{B}^{\times}/\widehat{R}^{\times}}\frac{\langle{f(g),h(g)}\rangle_{k-2\Sigma}}{w_{g}},$
@@
$w_{g}\in A^{\times},$
since $R_{g}^{1}:=\{x\in R_{g}^{\times}\ |\ {\rm N}(x)=1\}\subset R_{g,\text{tor}}^{\times}$ has prime factors dividing $q(q^{2}-1)$ and $R_{g}^{\times}/(O^{\times}\cdot R_{g}^{1})$ is a $2$-group.
Note that $\langle{x,H^{\prime}}\rangle_{k-2\Sigma}$ is non-zero modulo $\varpi$ if and only if there exists $H^{\prime\prime}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ with $H^{\prime\prime}(\widehat{B}^{\times})\subset A\rho(R_{\ell}^{\times})x$ such that
$v_{\ell}(\langle{H^{\prime\prime},H^{\prime}}\rangle)=0.$
@@
$\langle{H^{\prime\prime},H^{\prime}}\rangle=$ $\langle{H^{\prime\prime},\sum_{g\in R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}}\pi_{\mathfrak{q}}(g)H}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\langle{H^{\prime\prime},H}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\langle{H^{\prime\prime},\sum_{t\in K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})}\chi_{\mathfrak{q}}(t)\pi_{\mathfrak{q}}(t)\widehat{\varphi}}\rangle$
$=$ $\#R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}\cdot\gamma\#K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})\cdot\langle{H^{\prime\prime},\widehat{\varphi}}\rangle$.
Here the last equality follows from the multiplicity one of invariant pairings on $\pi$ (cf. [31, Lem. 2.6]) and the definition of $\gamma$.
Note that $R_{\mathfrak{q}}^{\times}/R_{\mathfrak{q}}^{\prime\times}$, $\gamma$ and $K_{\mathfrak{q}}^{\times}/(F_{\mathfrak{q}}^{\times}V_{\mathfrak{q}})$ are $\ell$-adic units as $\ell\nmid q(q^{2}-1)$. Recall that $\langle{\widehat{\varphi},x}\rangle_{k-2\Sigma}$ is non-zero modulo $\varpi$. Hence, the existence of $H^{\prime\prime}\in{\mathcal{M}}_{k}(\widehat{R}^{\times},{\mathrm{Frac}}(A))$ such that $H^{\prime\prime}(\widehat{B}^{\times})\subset A\rho(R_{\ell}^{\times})x$ and $\langle{\widehat{\varphi},H^{\prime\prime}}\rangle$ is non-zero modulo $\varpi$ follows, concluding the proof.